# ValveFit: An analysis-suitable B-spline-based surface fitting framework for patient-specific modeling of tricuspid valves

### v1

@@ 4.3 Fitting sensitivity with respect to noise
[…]
@@
00 $6.71\times 10^{-4}$ 0.0363 0.0118
$0.005$ $4.17\times 10^{-4}$ $0.0359$ $0.0118$
$0.01$ $6.96\times 10^{-4}$ $0.0352$ $0.0119$
$0.025$ $7.63\times 10^{-4}$ $0.0359$ $0.0122$
$0.05$ $9.61\times 10^{-4}$ $0.0376$ $0.0122$
$0.1$ $7.58\times 10^{-4}$ $0.0464$ $0.0138$
$0.2$ $9.25\times 10^{-4}$ $0.0590$ $0.0146$
Figure 11: Histogram showing the percentage of points with respect to sNND error at different point cloud noise levels (SD from $0.005$ to $0.2$) for simulation data. We observe that while the spread of errors slightly increases with increasing noise levels, the majority of the points lie in the lower error range, showing the robustness of ValveFit framework towards noise.
In Sections 4.14.3, we demonstrate the capability of the ValveFit framework to accurately fit point cloud data under different stages of valve motion, point cloud densities and Gaussian noise levels. When fitting point cloud data sequentially, the framework consistently maintains mean sNND error values between $0.0342$ and $0.0430$, even during complex valve motion. Table 2 shows that the ValveFit framework obtained a mean sNND value of $0.0155$ even with a relatively sparse point cloud density of only 50 points, which is comparable to the mean sNND values obtained at higher point cloud densities. As can be observed in Table 3, we also report a slight increase in sNND values as the amount of Gaussian noise in the point cloud increases. This validation study shows that the ValveFit framework exhibits fitting robustness under a range of conditions by accurately reconstructing surfaces that closely match the ground truth geometry across a range of noise levels and point cloud densities.
@@ 4.4 Surface Fitting for Patient Data
Figure 12: Patient 1 dataset: Fitted surfaces overlaid with the target point clouds at six stages of the cardiac cycle: (a) At the beginning of cardiac cycle $0.0\,T$, (b) $0.05\,T$, (c) $0.12\,T$, (d)$0.17\,T$, (e) $0.22\,T$, and (f) near the end-diastole $0.27\,T$. The first two rows show the top view of the fitted surfaces, while the bottom two rows show the side view of the fitted surfaces. Each point in the target point cloud is colored according to its sNND error.
In this section, we perform surface fitting using ValveFit framework on echocardiographic imaging data obtained from two patients. The point cloud data is labeled based on relative time, normalized with the duration of one cardiac cycle ($T$). In this normalized scale, $0.0\,T$ is defined as the beginning of the cardiac cycle, corresponding to the time of minimum right ventricular pressure. $1.0\,T$ represents the end of the same cardiac cycle and corresponds to the beginning of the next cardiac cycle ($1.0\,T_{i}=0.0\,T_{i+1}$). Important physiological events are also marked within a cardiac cycle, where $0.3\,T$ corresponds to end-diastole (ED), when the valve leaflets are coapted, and approximately at $0.45\,T$, which corresponds to end-isovolumetric contraction [77]. Figs. 12 and 13 show surface fitting results for the first patient data at $0.0\,T$, $0.05\,T$, $0.12\,T$, $0.17\,T$, $0.22\,T$, and $0.27\,T$), while Figs. 14 and 15 show the results for the second patient data at $0.75\,T$, $0.80\,T$, $0.85\,T$, $0.93\,T$, and $1.00\,T$).
@@
The template geometry shown in Fig. 2, is used for fitting patient data. The template surface parameterization consists of $13$ control points with degree $p=3$ basis functions along the $u$ direction and $5$ control points with degree $q=2$ basis functions along the $v$ direction. The weight parameters in Eq. (3) for the first patient data were set to $w_{\text{cd}}=20$, $w_{\text{hd}}=0.5$, $w_{\text{a}}=1$, $w_{\text{orth}}=10$, and $w_{\text{norm}}=5$. These weights are identical to the weights used for the validation study in Section 4.1, except that for fitting patient data we are using $w_{\text{a}}=1$ instead of zero. This is because we have access to point clouds where the annulus and leaflets have been separately segmented. For the second patient data, we observed higher noise intensity in the point clouds. To avoid overfitting the noise, we thus increased the relative importance of the regularization terms in the loss function (Eq. (2)) compared to the shape fidelity terms. Specifically, we decreased the weights $w_{\text{cd}}$ from $20$ to $5$ and $w_{\text{a}}$ from $1$ to $0.1$, ensuring that the surface does not try to overfit the noisy point cloud and remains smooth.
Figure 14: Patient 2 dataset: Fitted TV B-spline surfaces overlaid with the corresponding target point clouds captured at 5 stages of the cardiac cycle: (a) $0.75\,T$, (b) $0.80\,T$, (c) $0.85\,T$, (d) $0.93\,T$, and (e) at the end of the cardiac cycle ($1.00\,T$). The top two rows show the top view of the surfaces, while the bottom two rows show the side view. Each point in the target point cloud is colored according to its sNND error, providing a visual representation of the accuracy of the fitted surface at that location.
Figure 15: Patient 2 dataset: Percentage of points in the dataset plotted against the sNND error at different time points across the cardiac cycle. Each subplot corresponds to fitting accuracy at a certain time point in the cardiac cycle
Table 4: Patient 1 dataset: Minimum, maximum and mean sNND errors between the fitted B-spline surface and target point clouds at different stages of the cardiac cycle.
@@
$0.0\,T$ $9.35\times 10^{-6}$ $0.0904$ $0.0152$
$0.05\,T$ $4.35\times 10^{-6}$ $0.0599$ $0.0142$
$0.12\,T$ $5.47\times 10^{-6}$ $0.0629$ $0.0143$
$0.17\,T$ $1.84\times 10^{-5}$ $0.0744$ $0.0155$
$0.22\,T$ $7.82\times 10^{-6}$ $0.0982$ $0.0152$
$0.27\,T$ $5.33\times 10^{-6}$ $0.0896$ $0.0225$
@@
In Figs. 13 and 15, we show histograms of the sNND error distributions from the target point clouds to the fitted surface for Patients 1 and 2, respectively. For the Patient 1 dataset (Fig. 13), the maximum observed sNND error is $0.0982$, with the majority of points reporting sNND errors below $0.04$. The maximum observed sNND error for the Patient 2 dataset (Fig. 15) is $0.1359$, with a greater percentage of points reporting an sNND error below $0.06$. The distributions show consistent performance of the framework throughout the cardiac cycle. According to Table 4, the mean sNND errors for the Patient 1 dataset are between $0.0142$ and $0.0225$ across the six stages, showing the framework’s consistently stable performance throughout the cardiac cycle. The maximum sNND values ranged from $0.0599$ to $0.0982$, with slightly higher values at $0.22\,T$ and $0.27\,T$. These values were found to correlate with increased geometric complexity at the closed valve configurations.
Compared to the Patient 1 dataset, the sNND errors for the Patient 2 dataset (Table 5) were higher, with mean sNND values ranging from $0.0205$ to $0.0277$—approximately $40$ to $50\%$ greater than those observed in the Patient 1 dataset. Both the mean and maximum sNND errors for the Patient 2 dataset decrease from $0.75\,T$ to $1.0\,T$, as we fit sequentially from the closed to open valve configurations. The Patient 2 dataset required weight parameter adjustments as discussed above due to higher noise levels. The reduced relative importance of the Chamfer distance loss term compared to the regularization terms is the primary contributor to the increase in sNND error.
@@
Table 5: Patient 2 dataset: Minimum, maximum and mean sNND errors between the fitted B-spline surface and target point clouds at different stages of the cardiac cycle.
Step Minimum sNND Maximum sNND Mean sNND
$0.75\,T$ $4.14\times 10^{-5}$ $0.1359$ $0.0277$
$0.80\,T$ $4.59\times 10^{-6}$ $0.1273$ $0.0230$
$0.85\,T$ $6.41\times 10^{-6}$ $0.1241$ $0.0229$
$0.93\,T$ $1.40\times 10^{-5}$ $0.0924$ $0.0212$
$1.0\,T$ $3.86\times 10^{-4}$ $0.0931$ $0.0205$
The loss function in Eq. 2 does not contain any explicitly defined regularization terms to prevent global intersections, which may occur when fitting highly noisy or irregular point clouds. Intersections are avoided by employing the one-sided chamfer distance (Eq. 5) from the point cloud to the B-spline surface. Using the one-sided Chamfer distance prevents the issue of multiple surface points converging toward the same point in the point cloud when fitting extremely sparse point cloud data. Through hyperparameter tuning, the ValveFit framework balances geometric regularity and fitting accuracy when dealing with different noise levels. Thus, this framework can be used for generating analysis-suitable patient-specific B-spline surfaces for a variety of patient datasets with varying image quality and segmentation accuracy.
@@ 4.5 Tricuspid Valve Structural Simulation
Accurate modeling of TV closure under systolic loads is vital for predicting regurgitation and supporting surgical planning. To demonstrate the utility of our framework in patient-specific applications, we perform a simplified simulation on the fitted B-spline surface and highlight its ability to extract meaningful metrics such as strain. In this study, we adopt the isogeometric analysis framework developed by Johnson et al. [20] to simulate valve closure using a patient-specific surface fitted with ValveFit framework.
For this simulation, the point cloud for Patient 2 at $1.0\,T$ is considered to be the initial, stress-free configuration. The leaflet material behavior is modeled using the isotropic Lee–Sacks model [78], where the strain energy functional is defined as $\frac{1}{2}c_{0}(I_{1}-3)+\frac{1}{2}c_{1}(e^{c_{2}(I_{1}-3)^{2}}-1)$. Here, $I_{1}$ is the first invariant (trace) of the left Cauchy–Green deformation tensor, and $c_{0}$, $c_{1}$, and $c_{2}$ are the material coefficients set to $c_{0}$ = $10$ kPa, $c_{1}$ = $0.209$ kPa, and $c_{2}$ = $9.046$ [79]. The leaflet’s mass density is set at $1.0$ g/cm3 with a thickness of $0.0396$ cm.
@@
Accurate TV dynamics requires modeling the chordae tendineae, as healthy chordae play an important role in preventing valve prolapse. Without them, the valve cannot properly seal, leading to regurgitation and other functional impairments. Despite their importance to valve function, it is difficult for current non-invasive imaging techniques to capture chordae with enough fidelity for accurate geometric modeling. Previous efforts have worked around this issue by creating functional chordae maps [80] or using interactive dynamic simulations to estimate chordae location heuristically to match patient-valve behavior [81].
This work uses simplified chordae to prevent leaflet prolapse and demonstrate how the geometry is directly analysis suitable through closure mechanics. Six idealized cables, each 1 cm in length, are attached to the leaflet and extend outwards perpendicular to its surface. Each TV leaflet has two chordae, with one end of each chordae attached to an insertion point along the free edge of the TV and the other fixed in space to represent papillary muscle attachment. An IGA cable formulation [20] is used for the chordae, which are modeled as cubic B-spline curves with Young’s modulus of $40$ MPa and a fictitious radius of $0.023$ cm. This setup ensures that the valve is constrained and resists leaflet prolapse. For more accurate patient-specific simulations, proper insertion points and papillary muscle locations can be used for more clinically viable results.
The simulations specifically focus on the tricuspid valve in its closed state. For simplification, the positions of the annulus and papillary muscles are fixed in the initial configuration of $1.0\,T$, and thus their motion or deformation is not modeled. Valve closure is achieved by applying a pressure load on the ventricular side of the leaflet surface. This load increases linearly from 00 to $25$ mmHg over the course of $5$ milliseconds (00 to $0.005$ s), mirroring the physiological pressure gradients across the TV during systole.
Figure 16 shows the top and side views of the simulated valve closure, highlighting the framework’s ability to produce patient-specific, analysis-suitable geometries for predictive modeling. These geometries allow for the recovery of valuable cardiac cycle metrics, including deformation, coaptation, and strain (e.g. MIPE, the maximum in-plane principal Green–Lagrange strain), which are often inaccessible through clinical imaging alone. With access to more complete patient data, it should be possible to recover physiologically relevant, patient-specific metrics throughout the entire cardiac cycle.

### v2

@@ 5.3 Fitting Sensitivity with Respect to Noise
In Fig. 15, we evaluate the robustness of the ValveFit framework under varying levels of noise in the point cloud data. Here, random perturbations were added to the point cloud consisting of $2500$ points at simulation step $60$. The goal was to simulate the noise typically observed in point cloud data derived from echocardiographic images (Section 3). Gaussian noise was added with increasing standard deviations (SD) of $0.005,0.01,0.025,0.05,0.1,\text{and }0.2$ [90], representing a specific degree of perturbation. Despite increasing noise, the ValveFit framework successfully captures the finer features of the surface while simultaneously generating smooth and intersection-free surfaces. Even while the noise levels increase, the fitted surfaces maintain a close approximation of the underlying geometry without overfitting. The mean sNND values gradually increase from $0.0039$ to $0.0071$ as the noise levels rise, according to the sNND errors in Table 3. In Fig. 16, we similarly observe a few points falling into high error bins as the noise levels rise. However, the sNND error does not significantly increase, showcasing minimal distortion of the B-spline surfaces due to noise.
@@
Figure 16: Histogram showing the percentage of points with respect to sNND error at different point cloud noise levels (SD from $0.005$ to $0.2$) for simulation data. We observe that while the spread of errors slightly increases with increasing noise levels, the majority of the points lie in the lower error range.
@@
$0.005$ $1.48\times 10^{-4}$ $8.90\times 10^{-3}$ $3.86\times 10^{-3}$
$0.01$ $3.90\times 10^{-4}$ $9.45\times 10^{-3}$ $3.90\times 10^{-3}$
$0.025$ $2.53\times 10^{-4}$ $1.12\times 10^{-2}$ $3.99\times 10^{-3}$
$0.05$ $4.67\times 10^{-4}$ $1.09\times 10^{-2}$ $4.35\times 10^{-3}$
$0.1$ $2.99\times 10^{-4}$ $1.95\times 10^{-2}$ $5.10\times 10^{-3}$
$0.2$ $3.81\times 10^{-4}$ $2.33\times 10^{-2}$ $7.14\times 10^{-3}$
In Sections 5.15.3, we demonstrate the capability of the ValveFit framework to accurately fit point cloud data under different stages of valve motion, point cloud densities, and Gaussian noise levels. When fitting point cloud data sequentially, the framework consistently maintains mean sNND error values between $0.0025$ and $0.0054$, even during complex valve motion. Table 2 shows that the ValveFit framework obtained a mean sNND value of $0.0032$, even with a relatively sparse point cloud density of only $250$ points, which is comparable to the mean sNND values obtained at higher point cloud densities. As can be observed in Table 3, we also report a gradual increase in sNND values as the amount of Gaussian noise in the point cloud increases. This validation study shows that the ValveFit framework exhibits fitting robustness under a range of conditions by accurately reconstructing surfaces that closely match the ground truth geometry across various noise levels and point cloud densities.
@@ 5.4 Surface Fitting for Patient Data
After validating the VALVEFIT framework in Sections 5.1-5.3, we demonstrate its performance for point clouds obtained from echocardiographic imaging data. The point cloud datasets are labeled based on relative time, normalized with the duration of one cardiac cycle ($T$). In this normalized scale, 0.0T0.0\,T is defined as the beginning of the cardiac cycle, corresponding to the time of minimum right ventricular pressure. 1.0T1.0\,T represents the end of the same cardiac cycle and corresponds to the beginning of the next cardiac cycle ($1.0\,T_{i}=0.0\,T_{i+1}$). Important physiological events are also marked within a cardiac cycle, where 0.3T0.3\,T corresponds to end-diastole (ED), when the valve leaflets are coapted, and approximately at 0.45T0.45\,T, which corresponds to end-isovolumetric contraction [91]. Figs. 17 and 18 show surface fitting results for the first patient data at 0.0T0.0\,T, 0.05T0.05\,T, 0.17T0.17\,T, 0.22T0.22\,T, 0.27T0.27\,T, and 0.34T0.34\,T, while Figs. 19 and 20 show the results for the second patient data at 0.93T0.93\,T, 1.0T1.0\,T, 1.05T1.05\,T, 1.10T1.10\,T, 1.18T1.18\,T, and 1.23T1.23\,T.
Figure 17: Patient 1 dataset: Fitted surfaces overlaid with the target point clouds at six stages of the cardiac cycle: (a) At the beginning of cardiac cycle 0.0T0.0\,T, (b) 0.05T0.05\,T, (c) 0.17T0.17\,T, (d) 0.22T0.22\,T, (e) 0.27T0.27\,T, and (f) and 0.34T0.34\,T. The first two rows show the side view of the fitted surfaces, while the bottom two rows show the top view of the fitted surfaces. Each point in the target point cloud is colored according to the sNND error.
@@
The template geometry shown in Fig. 2 is used for fitting patient data. The template surface parameterization for fitting the Patient 1 dataset consists of $23$ control points with degree $p=3$ basis functions along the $u$ direction and $6$ control points with degree $q=3$ basis functions along the $v$ direction, resulting in a control grid of size $23\times 6$. The weight parameters in Eq. (3) for the first patient data were set to $w_{\text{CD}}=20$, $w_{\text{HD}}=0.5$, $w_{\text{a}}=1$, $w_{\text{orth}}=10$, $w_{\text{TPE}}=1.5$, and $w_{\text{norm}}=5$. The weights are identical to those used in Section 5.1, except for an increased Tangent Point Energy weight ($w_{\text{TPE}}$) to account for larger inter-point distances. Additionally, for patient data, $w_{\text{a}}$ is set to a non-zero value because we have access to point clouds where the annulus and leaflets have been separately segmented. For the Patient 2 dataset, we observed greater geometric complexity and higher noise intensity in the point clouds. The template surface parameterization is changed to result in a control grid of size $23\times 8$. To avoid overfitting to the noise, we thus increased the relative importance of the regularization terms in the loss function (Eq. (2)) by increasing the weights $w_{\text{orth}}$ to $15$ and $w_{\text{norm}}$ to $7.5$, ensuring that the surface does not attempt to overfit the noisy point cloud and remains smooth.
Figure 19: Patient 2 dataset: Fitted TV B-spline surfaces overlaid with the corresponding target point clouds captured at 5 stages of the cardiac cycle: (a) 0.93T0.93\,T, (b) 1.0T1.0\,T, (c) 1.05T1.05\,T, (d) 1.10T1.10\,T, 1.18T1.18\,T, and (e) at the end of the cardiac cycle (1.23T1.23\,T). The first two rows show the top view of the surfaces, while the last two rows show the side view. Each point in the target point cloud is colored according to its sNND error, providing a visual representation of the accuracy of the fitted surface at that location.
Figure 20: Patient 2 dataset: Percentage of points in the dataset plotted against the sNND error at different time points across the cardiac cycle. Each subplot corresponds to fitting accuracy at a certain time point in the cardiac cycle.
Table 4: Patient 1 dataset: Minimum, maximum and mean sNND errors between the fitted B-spline surface and target point clouds at different timesteps.
@@
0.0T0.0\,T $1.2566\times 10^{-6}$ 0.04120 0.00995
0.05T0.05\,T $1.2560\times 10^{-6}$ 0.05581 0.00881
0.17T0.17\,T $2.8330\times 10^{-6}$ 0.05728 0.00871
0.22T0.22\,T $3.0224\times 10^{-6}$ 0.05310 0.01019
0.27T0.27\,T $1.6789\times 10^{-6}$ 0.05436 0.01357
0.34T0.34\,T $1.0263\times 10^{-6}$ 0.05481 0.01354
@@
In Figs. 18 and 20, we show histograms of the sNND error distributions from the target point clouds to the fitted surface for Patients 1 and 2, respectively. For the Patient 1 dataset (Fig. 18), the maximum observed sNND error is $0.05728$, with the majority of points reporting sNND errors below $0.03$. The maximum observed sNND error for the Patient 2 dataset (Fig. 20) is $0.07187$, with a greater percentage of points reporting an sNND error below $0.03$. The distributions show consistent performance of the framework throughout the cardiac cycle. According to Table 4, the mean sNND errors for the Patient 1 dataset are between $0.00995$ and $0.01357$ across the six stages, showing the framework’s consistently stable performance throughout the cardiac cycle. The maximum sNND values ranged from $0.0412$ to $0.05728$, with slightly higher values at time points 0.17T0.17\,T and 0.34T0.34\,T. These values were found to correlate with increased geometric complexity and leaflet coaptation, respectively.
@@
Table 5: Patient 2 dataset: Minimum, maximum and mean sNND errors between the fitted B-spline surface and target point clouds at different timesteps.
Timestep Minimum sNND Maximum sNND Mean sNND
0.93T0.93\,T $4.1777\times 10^{-6}$ 0.07178 0.01082
1.0T1.0\,T $1.9923\times 10^{-6}$ 0.03468 0.00976
1.05T1.05\,T $2.1534\times 10^{-6}$ 0.04140 0.00947
1.10T1.10\,T $4.5950\times 10^{-6}$ 0.05571 0.01082
1.18T1.18\,T $2.2556\times 10^{-6}$ 0.06232 0.01262
1.23T1.23\,T $1.1482\times 10^{-6}$ 0.07187 0.01593
Compared to the Patient 1 dataset, the sNND errors for the Patient 2 dataset (Table 5) were slightly higher, with mean sNND values ranging from $0.00947$ to $0.01593$. The fitting accuracy varies throughout the cardiac cycle. Both the mean and maximum sNND errors decrease initially, reaching a minimum value near the fully open valve configuration around 1.0T1.0\,T and 1.05T1.05\,T. Subsequently, the errors trend upward through the remainder of the cycle. The Patient 2 dataset required weight parameter adjustments, as discussed above, due to higher noise levels in the point clouds. The reduced relative importance of the Chamfer distance loss term compared to the regularization terms is the primary contributor to the slight increase in sNND error. Through hyperparameter tuning, the ValveFit framework balances geometric regularity and fitting accuracy while dealing with different noise levels. Thus, this framework can be used to generate analysis-suitable patient-specific B-spline surfaces for a variety of patient datasets with varying image quality and segmentation accuracy. In Fig. 21, we present the fitted B-spline surfaces overlaid on echocardiographic slices, together with the target points segmented from the corresponding slices for the representative time points of Patients 1 and 2 datasets. The slices are extracted along planes oriented orthogonally to either the x-axis (sagittal) or y-axis (coronal), with the slicing positions specified by the coordinates shown above each panel. The overlays demonstrate a close correspondence between the fitted surfaces and the key anatomical structures in the echocardiographic slices, providing visual confirmation of the framework’s accuracy. These results highlight the capability of the proposed approach to generate patient-specific, analysis-suitable surface representations that are consistent with clinical imaging data.
Figure 21: Overlays of the fitted B-spline surfaces (blue) and segmented target point clouds (red) on the corresponding 2D echocardiography slices for Patient 1 at (a) 0.0 T and (b) 0.32 T, and for Patient 2 at (c) 1.05 T and (d) 1.32 T. Each panel demonstrates the agreement of the fitted patient-specific surfaces with the underlying image data.
@@ 5.5 Tricuspid Valve Structural Simulation
Accurate modeling of TV closure under systolic loads is vital for predicting regurgitation and supporting surgical planning. To demonstrate the utility of our framework for generating analysis-suitable B-spline surfaces in patient-specific applications, we perform a simplified structural simulation on the fitted B-spline surface and highlight its ability to extract meaningful metrics, such as strain. For the atrioventricular heart valves, the primary driving force for valve opening and closure is the pressure gradient across the valve leaflet surfaces, which can be idealized or realized by means of the pressure loading in the structural mechanics analysis [92, 93, 20, 94, 95, 13, 23, 96, 40, 97, 41, 98, 39]. In this study, we adopt the isogeometric analysis framework in  [20] to simulate valve closure using a patient-specific surface fitted with ValveFit framework.
For this simulation, the point cloud for Patient 2 at 1.0T1.0\,T is considered the initial, stress-free configuration. The leaflet material behavior is modeled using the isotropic Lee–Sacks model [99], where the strain energy functional is defined as $\frac{1}{2}c_{0}(I_{1}-3)+\frac{1}{2}c_{1}(e^{c_{2}(I_{1}-3)^{2}}-1)$. Here, $I_{1}$ is the first invariant (trace) of the left Cauchy–Green deformation tensor, and $c_{0}$, $c_{1}$, and $c_{2}$ are the material coefficients set to $c_{0}$ = $10$ kPa, $c_{1}$ = $0.209$ kPa, and $c_{2}$ = $9.046$ [100]. The leaflet’s mass density is set at $1.0$ g/cm3 with a thickness of $0.0396$ cm.
@@
Accurate TV dynamics require modeling the chordae tendineae, as healthy chordae play an important role in preventing valve prolapse. Without them, the valve cannot properly seal, leading to regurgitation and other functional impairments. Despite their importance to valve function, it is difficult for current non-invasive imaging techniques to capture chordae with enough fidelity for accurate geometric modeling. Previous efforts have worked around this issue by creating functional chordae maps [101] or by using interactive dynamic simulations to estimate chordae locations heuristically to match patient-valve behavior [102].
This work uses simplified chordae to prevent leaflet prolapse and demonstrates how the geometry is directly analyzable through closure mechanics. Six idealized cables, each 1 cm in length, are attached to the leaflet and extend outward, perpendicular to its surface. Each TV leaflet has two chordae, with one end of each chordae attached to an insertion point along the free edge of the TV, and the other fixed in space to represent papillary muscle attachment. An IGA cable formulation [20] is used for the chordae, which are modeled as cubic B-spline curves with a Young’s modulus of $40$ MPa and a fictitious radius of $0.023$ cm. This setup ensures that the valve is constrained and resists leaflet prolapse. For more accurate patient-specific simulations, proper insertion points and papillary muscle locations can be used to achieve more clinically viable results.
The simulations specifically focus on the tricuspid valve in its closed state. For simplification, the positions of the annulus and papillary muscles are fixed in the initial configuration of 1.0T1.0\,T, and thus their motion or deformation is not modeled. Valve closure is achieved by applying a pressure load on the ventricular side of the leaflet surface. This load increases linearly from 0 to $25$ mmHg over the course of $5$ milliseconds (0 to $0.005$ s), mirroring the physiological pressure gradients across the TV during systole.
Fig. 22 shows the top and side views of the simulated valve closure, highlighting the framework’s ability to produce patient-specific, analysis-suitable geometries for predictive modeling. These geometries allow for the recovery of valuable cardiac cycle metrics, including deformation, coaptation, and strain (e.g. MIPE, the maximum in-plane principal Green–Lagrange strain), which are often inaccessible through clinical imaging alone. With access to more complete patient data, it should be possible to recover physiologically relevant, patient-specific metrics throughout the entire cardiac cycle.